OpenStudy (anonymous):

What is the sum of the arithmetic sequence 6, 11, 16 …, if there are 24 terms?

OpenStudy (anonymous):

the common difference is 5,first term is 6..so at first find the 24th term..

OpenStudy (anonymous):

There are 2 approaches: first, you can simply use the formula for arithmetic series (i.e. sum of the sequence) ---- sum = 24 (a1 + a24)/2... a1 is 6 and, a24 is 23(5) + 6 = 121...

OpenStudy (anonymous):

But, if you want to view it intuitively, consider this: the first term is 6 and the last term is 121... 6 + 121 = 127... the second term is 11 and the 23rd term is 116... 11 + 116 = 127... so, every pair is 127.... And, there are 12 pairs... 12 pairs x 127 = 1524